A company uses the formula C(x)= 0.02x^2+3.4x+150 to model the unit cost in dollars for producing x stabilizer bars. For what number of bars is the unit cost at it's minimum? What is the unit cost at that level of production?
Answer by stanbon(75887)

A company uses the formula C(x)= 0.02x^2+3.4x+150 to model the unit cost in dollars for producing x stabilizer bars. For what number of bars is the unit cost at it's minimum? What is the unit cost at that level of production?
------------------
The equation is a quadratic with a=0.02 , b=3.4 , c=150
The minimum is at x=-b/2a = -3.4/(2*0.02)= -85
------------
Comment:
The number of bars cannot be negative. Check you equation again.
================
Cheers,
Stan H.